# Concept Formation and Alignment in Language Models: Bridging Statistical Patterns in Latent Space to Concept Taxonomy

### v1

@@ 3.2 Louvain Community Detection
Then, it aggregates the communities to identify hierarchical structures. In this phase, each community is considered as a single node and the links between the new nodes are calculated as the sum of the weight of the links between nodes in the corresponding two communities. More details are given in Algorithm 1.
@@ 15       Reduce each community to a single node;
3.3 Concept Extraction Algorithm
for our concept extraction algorithm, as the first step, we create and weight the adjacency graph using $K$-NN, UMAP-based weighing formula (mentioned in section 3.1), and then use Louvain [6] and label propagation [58] community detection algorithms.
There are multiple viable strategies to extract hierarchical communities in our methodology; the first one is to use algorithms such as Louvain that inherently generate hierarchical communities [6]. However, the summation of the weights as a new weight in the community aggregation phase of Louvain algorithm skews the weighted graph in favor of merging smaller communities in the next phase. This detaches the community detection from the actual values in the semantic representation space (i.e. graph weights higher in the Louvain hierarchy no longer reflect the geometrical affinity of the nodes). Thus, for our concept extraction algorithm, we use 1-2 community aggregation and rather use KNN iteratively with different granularity for extracting hierarchical concepts (more details are described in Algorithm 2).
Note that we only use well-established methods such as $K$-NN, UMAP-based weighing formula, as well as Louvain [6] and label propagation [58] community detection algorithms to capitalize on the established generalizability of these algorithms. Although, as we show in the next sections that our method produces amazingly good categories, it should be noted that we intend to focus our analysis on “if the language model forms concepts" rather than creating the most optimal concept extraction mechanism. Thus, as an extension to this work, one can focus on further optimizing our proposed method. Notably, our concept extraction is algorithm-agnostic; alternatives could be readily employed.
Data: All tokens in the input embeddings.
Result: A set of hierarchical communities.
1 Create a community list. The initial community is the entire space;
2 for k=[different neighbor sizes] do
3       for all communities do
4             (1) Generate KNN graph of the community using the input embedding weights;
5             (2) Compute the edge weights of the graph using fuzzy simplex;
6             (3) Apply Louvain/Label-Propagation community detection;
7             (4) Add the identified communities to the list;
8
9
Algorithm 2 Concept Extraction
3.4 Models Under Investigation
Using our proposed method, we investigated the embedding spaces of GloVe, ALBERT, and T5 to determine if concept formation occurs consistently across these diverse language models. We chose this spectrum of models to: (i) assess if our method for identifying concept formation is applicable across a wide range of architectures and training data, and (2) investigate whether the concept formation is a fundamental property of language representation, regardless of the specific model architecture or training data. We hypothesize that the language model with the least embedding layer dimension is more appropriate for our investigative analysis, because, when all other factors are held constant, models with more compact latent spaces generally produce more invariant representations [20]. Thus, for the detailed analysis and evaluations, we pick ALBERT  [^2] We used open source Hugging Face albert-base-v2 and albert-xxlarge-v2 (https://huggingface.co/albert). [23] as our primary target since it has the tightest latent space dimensionality.
Finally, we hypothesize that exploring the formation of such analogs early in the language model, potentially at the input embedding layer, could open the door to the possibility of knowledge-reasoning separation. Thus, we focus our investigation on this layer. More details about the models can be found in Appendix A.
4 Experiments and Discussions
@@
Concept formation involves the creation of abstract internal representations through the clustering of inputs with shared features. While numerous intrinsic clustering metrics exist to assess cluster formation quality, their application to our use case is limited by two factors:
@@
  • The formed clusters are not situated within Euclidean space [24, 8, 7], rendering geometric properties such as cluster distances inadequate indicators of concept well-formedness and distinctiveness.
@@
  • Concepts inherently possess a degree of vagueness [19], thus metrics like compactness or separation do not reliably reflect the quality of formed concepts.
Consequently, to assess the quality of the formed concepts, we shift our focus to evaluating concept alignment, specifically examining both LM-LM and Human-LM alignment. LM-LM alignment allows for the generalization of our observations to other language models, while Human-LM alignment serves as an indicator of the quality and real-world interpretability of the formed concepts.
Concept Extraction Setup. We configured our algorithms to create $k$-NN graph iteratively for different values of $k$. This enables us to observe the communities/concepts formation at various granularities. Table 1 shows the number of identified clusters for $k=[100,75,50,25,12,6]$.
Table 1: Number of communities with different granularities of nearest neighbors for ALBER, T5 and Glove. For Glove, we only used the subset of GloVe that present in Albert Vocab
$K$ settings Vocab
Models 100 75 50 25 12 6 Size
T5 1 115 1203 4551 8137 9407 32000
Albert 8 133 1058 4442 7718 8626 30000
GloVe 9 207 1157 3521 6237 7200 25869
4.1 LM-LM alignment
To evaluate LM-LM alignment, we measure the community mapping across the language models under test. We define a mapping as the ratio of the successfully mapped communities from the embedding space of the language model with the smaller vocabulary to the language model with the larger vocabulary. Furthermore, we define success in mapping process as if the ratio of of found tokens from source community in the destination community is greater than a certain threshold. thresholds greater than 0.5 imply at worst many-to-one mapping (when the number of tokens in the source and destination space are similar, it is more likely to imply one-to-one mapping). Such one-to-one mapping implies a direct concept alignment, and many-to-one mapping implies the source community is aligned with hierarchies of the destination concept.
This mapping score provides a quantitative measure of the degree to which the concept communities formed by one model can be translated to the other. To refine our analysis, we filtered the tokens to include only those that start with whitespace in T5 and ALBERT models. Empirically, this enhances the likelihood of associating the token with a word existing in the external world, thereby enhancing the validity of our concept alignment assessment. Note that to find the mappings we only used the token sets that are shared between the model pairs. That is 25869 for Albert/GloVe, 14861 for T5/GloVe and 5055 tokens for T5/Albert pairs.
Table 2: Pair-wise Mapping score of the formed communities in the Glove, Albert, and T5 embedding spaces.
Thre- Albert / GloVe T5 / GloVe T5 / Albert
shold $K$=75 $K$=50 $K$=25 $K$=12 $K$=6 $K$=75 $K$=50 $K$=25 $K$=12 $K$=6 $K$=75 $K$=50 $K$=25 $K$=12 $K$=6
0.5 0.70 0.57 0.62 0.68 0.70 0.62 0.74 0.84 0.89 0.91 0.61 0.78 0.91 0.95 0.97
0.6 0.56 0.42 0.43 0.42 0.42 0.46 0.60 0.70 0.72 0.72 0.43 0.60 0.80 0.86 0.86
0.7 0.42 0.30 0.30 0.28 0.27 0.27 0.46 0.54 0.56 0.56 0.30 0.49 0.70 0.79 0.80
0.8 0.32 0.22 0.24 0.23 0.21 0.15 0.39 0.48 0.52 0.52 0.24 0.45 0.69 0.78 0.79
0.9 0.17 0.16 0.20 0.21 0.21 0.07 0.32 0.45 0.50 0.51 0.18 0.43 0.67 0.78 0.79
[…]
The aforementioned observations strongly support the generalization of LM-LM concept alignment. The high degree of alignment, particularly for granular concepts and between models with similar architectures (like ALBERT and T5), indicates a shared understanding of semantic relationships across these models. This opens up exciting possibilities for leveraging this alignment to improve model performance or transfer knowledge between models without the need for extensive retraining. However, further research is needed to explore the extent and limitations of this potential knowledge transfer, especially for more abstract concepts and between models with different architectures.

### v3

@@ 3.2 Louvain Community Detection
Then, it aggregates the communities to identify possible hierarchical structures. In this phase, each community is considered as a single node and the links between the new nodes are calculated as the sum of the weight of the links between nodes in the corresponding two communities. More details are given in appendix C, algorithm 2.
3.3 Concept Extraction Algorithm
For our concept extraction algorithm, as the first step, we create and weight the adjacency graph using $K$-NN, UMAP-based weighing formula (mentioned in section 3.1), and then use Louvain algorithm. Algorithm 1 describes the concept extraction process (see appendix E for details on the algorithm methodology and considerations). We configured our algorithms to create $k$-NN graph iteratively for different values of $k$. This enables us to observe the communities/concepts formation at various granularities. Table 1 shows the number of identified clusters for $k=[100,75,50,25,12,6]$.
Table 1: Number of communities with different granularities of nearest neighbors for Albert, T5, Glove, and Llama3. For Glove, we only used the subset of GloVe that present in Albert vocabulary
$K$ settings Vocab
Models 100 75 50 25 12 6 Size
T5 1 115 1203 4551 8137 9407 32000
Albert 8 133 1058 4442 7718 8626 30000
GloVe 9 207 1157 3521 6237 7200 25869
Llama-3 7 23 844 6044 18644 32535 128256
Hierarchy Formation. When examining the broader perspective (i.e., $k$=100), the model primarily found groups of named entities (names of people and places), adverbs, sub-words, some number symbols, and etc. (appendix J, Figure 5 shows the overall concept hierarchies of the Albert vocab). Zooming in further (e.g., $k$=75), these communities revealed more specific clusters that are relatable to the real world. For instance, within named entities, clusters formed for personal vs. location names, even further pinpointing locations by country. As the granularity level increases (approaching a smaller $k$ value), clusters exhibit a stronger association with word forms.
Figure 1: The identified name and location communities for different k granularity for Albert model. At the left-side the average precision score for the extracted graph within each granularity is given. For the more detailed tables, see H (Note that the results for other models are also available in appendix I).
Figure 2: Visualization of the identified name and location communities of size larger than 10 entities. UMAP projection along with Seaborn 57 is used for the visualization.
4 Evaluation: Alignment with External Knowledge
We extracted several meaningful conceptual communities, ranging from symbo-numerical groups, to concrete objects and Named-Entities (such as plant groups, animals, car brands, names, and locations), to more abstract groupings such as colors, social roles and structures, currencies, etc. (more details in appendix J).
To evaluate RQ1 and RQ2, we selected named entities as well as numbers, primarily because the existence of external evaluation datasets for NE as well as the well defined nature of numbers allow us to examine (1) the alignment with external world knowledge, and (2) to observe within group properties of the conceptual communities. To assess the consistency of conceptual groupings across different LLMs, we applied our method to the embedding spaces of GloVe 40, Albert 23, T5 42, and Llama-3 1 (See Appendix F). For brevity, we present the overview of the results for Albert in the main section (for more detailed table for the all aforementioned models please refer to Appendix H and I)
4.1 Named Entities
To evaluate the quality and the alignment of the extracted conceptual grouping with the external world, we used we leveraged the name-dataset 43 and the country-state-city database 12. These datasets provide comprehensive information on names (including gender and country-specific popularity) and locations (including hierarchical relationships between countries, states, and cities), enabling a rigorous assessment of the conceptual groupings (see appendix H for more details).
The granularity of observed clusters was influenced by both the number of constituent samples and conceptual attributes. For instance, English/American named entities and locations tended to form distinct clusters at higher levels of granularity (k values) due to their high frequency within the vocabulary. However, upon further increasing granularity (higher k values), hierarchical structure emerged within these clusters. The conceptual attributes contributing to this hierarchical organization within named entities included (1) entity type (human name or location), (2) name part (first or last), gender, and (3) regional/national origin.
Note that since there’s no one-to-one mapping between names/locations in LLM vocabulary and external datasets (e.g., a name may appear in multiple countries, the external dataset is also a superset), recall is less relevant. We prioritize precision to evaluate accuracy, as it better reflects our ability to identify correct matches. We observe high precision score across the identifies name entity cluster. Figure, 1 shows the main groupings for Albert Model. We got average precision score of 0.81 for K=75 and 0.76 for K=50/25. Appendix H.1 contains the detailed tables for different models. Furthermore, we observed a degree of geographic ordering within the identified communities. As illustrated in Figure 2, there appears to be a general trend from east to west as we move across the communities from bottom-left to top-right. The leftmost communities are predominantly associated with Japanese locations and names, while those on the rightmost side are primarily linked to Europe and the United States. This suggests that the model’s internal representations in the input embedding layer may inherently capture geographical relationships.
Table 2: Topological ordering score of different communities of the numbers. Years refer to the cluster of numbers between 1816 - 2021. Support refers to the number of samples within the cluster.
Category k-1 k-3 k-5 Support
0-100 0.68 0.89 0.92 100
100-262 0.75 0.94 1.0 115
263-300 0.46 0.78 0.89 37
300-400 0.42 0.71 0.84 82
400+ 0.53 0.75 0.83 110
Years 0.86 0.96 0.98 203
4.2 Symbols-Numbers
At a high level, symbols and numbers are distinctly separated from other tokens in the embedding space. Within their own domain they form communities according to (i) years, (ii) integer values, (iii) tokens indicating monetary values (e.g., \$1), (iv) ratio/time (e.g., 2:1, 3:30), (v) fractions (e.g., ’/8’), (vi) large values (e.g., ‘100,000’), and (vii) percentages (e.g., $42\%$). In these communities, integers create sub-communities based on their hundreds.
To align with the human notion of numbers, it is essential to identify a topological ordering within the numerical embeddings. As the embedding manifold may not conform to a Euclidean structure 24, 9, 8, conventional distance-based order measures were inapplicable. Lemma 5 formulates the local ordering of embeddings within a embedding manifold.
Lemma 1.
Local Ordering on a Manifold: Let M be a manifold and let d(a, b) denote the distance between points a and b on M. For a given positive integer $k$, we say that a point x on M is locally ordered if and only if:
$x\in top_{k}(x+1)\cap top_{k}(x-1)$ (5)
where $top_{k}(a)$ denotes the set of k-nearest neighbors of point a on M.
This localized approach captures the intuitive concept of topological ordering within a small neighborhood, rather than relying on a global ranking.
Topological Ordering Score is defined as:
$S=\frac{1}{n}\sum\limits_{i=0}^{n}{f_{k}(x_{i})}$ (6)
where $n$ indicates the number of embeddings in the given cluster, $k$ is an integer that controls the strictness of ordering, and $f_{k}(x_{i})$ is boolean function that returns 1 if $x_{i}$ hold lemma 5 condition.
Table 2 presents the topological ordering scores of the communities identified by our algorithm. The "k-1" score represents strict topological ordering, where $x+1$ lies exactly between $x$ and $x+2$ in the embedding space. The "k-3" and "k-5" scores measure relaxed topological ordering, where $x+1$ falls within the 3 or 5 nearest neighbors of both $x$ and $x+2$, respectively. As the results indicate, all detected communities exhibit a high degree of local topological ordering, regardless of the chosen level of strictness. This finding is significant because it suggests that LLMs may possess not only the ability to categorize heterogeneous input entities but also the capacity to construct meaning within smaller, internally consistent structures (internally homogeneous sub-structures). This implies the potential for LLMs to move beyond simply classifying information to actively interpreting and generating meaning within specific clusters.
5 LLM-LLM alignment
Investigating LLM-LLM alignment in the input embedding layer is crucial to answer RQ3, because it reveals how well different LLMs represent the same concepts in their initial processing stages. This analysis provides valuable insights into the influence of model architecture, size, and training regimes on the formation of language representations. To quantify this alignment, we defined an alignment score that specifically captures the discrete overlaps of nearest neighbors in embedding spaces. This score offers a precise assessment of how similarly the LLMs represent shared tokens, enabling a more nuanced understanding of their representational alignment.
We define Alignment Score as:
$S=\frac{1}{n}\sum\limits_{j=0}^{j=n}\frac{|top\_k({LLM}^{1}_{j})\cap top\_k({LLM}^{2}_{j})|}{k}$ (7)
where ${LLM}^{1}$ and ${LLM}^{2}$ denote the two LLMs under investigation, $n$ represents the number of shared tokens between the LLM pair, and $top\_k({LLM}^{1}_{j})$ and $top\_k({LLM}^{2}_{j})$ represent the sets of the top k nearest neighbors of the $j$-th token in ${LLM}^{1}$ and ${LLM}^{2}$ respectively. The alignment score for two randomly chosen embeddings with the same vocabulary is 0 (assuming $k<<n$).
We calculated this alignment score across a range of LLMs, encompassing diverse architectures and training objectives. we first identify the set of tokens shared across the vocabularies of a given pair of LLMs. For each shared token, we compute its k nearest neighbors ($k=3,5,10,50$) in the embedding space of both models. We use pairwise cosine similarity measure to find top-k tokens within an LLM. This selection included Albert (encoder-based), T5 (encoder-decoder), Llama 70B (decoder-only) to examine the impact of high-level model architecture. To investigate the effect of model size, we included T5 models of different scales: T5-small, T5-base, T5-large, T5-3B, and T5-11B.
Table 3: Pair-wise alignment score for various LLM sizes, and training regimes. higher alignment score is better.
Model-1 Model-2 k=3 k=5 k=10 k=50 Support
Albert-XXL Albert-base 0.501 0.490 0.472 0.438 30000
Albert-XXL Albert-L 0.56 0.54 0.52 0.48 30000
Albert-XXL Albert-XL 0.60 0.58 0.55 0.51 30000
T5-11B T5-small 0.61 0.577 0.52 0.38 32100
T5-11B T5-base 0.68 0.66 0.60 0.51 32100
T5-11B T5-large 0.74 0.71 0.66 0.61 32100
T5-11B T5-3B 0.79 0.76 0.71 0.63 32100
Llama3-70B Llama3-1B 0.57 0.55 0.50 0.31 128256
Llama3-70B Llama3-3B 0.56 0.53 0.48 0.27 128256
Llama3-70B Llama2-70B 0.65 0.61 0.56 0.47 22430
Table 4: Cross architecture alignment score for various LLM architectures, sizes, and training regimes. higher alignment score is better. Support column indicates the number of shared tokens between the LLM pair.
Model-1 Model-2 k=3 k=5 k=10 k=50 Support
Albert-XXL Llama3-3B 0.55 0.52 0.49 0.44 18640
Albert-XXL T5-3B 0.63 0.61 0.59 0.54 12307
Llama3-3B T5-3B 0.66 0.62 0.58 0.53 20603
Llama3-70B T5-11B 0.64 0.61 0.56 0.47 20603
Analyzing results from tables 3 and 4, we observe:
  • Table 3 shows that model size is a contributing factor in achieving higher alignment scores when comparing within the same architecture. Since larger models also demonstrate better performance on benchmark tasks, we can infer that the quality of the concepts formed in the input embedding layer is positively correlated with model size, assuming architecture and training regimes are held constant.
  • Both tables 3 and 4 show that alignment scores generally decrease as the value of k increases. This trend suggests that while models share a core understanding of semantic similarity at smaller k values, they diverge when considering more generic concepts at larger k values. The alignment drop is larger in decoder-only models, likely due to their unidirectional context  [^2] This is consistent with human studies showing that readers’ eye movements are bidirectional: forward to absorb new information, and backward to resolve comprehension issues or correct errors 48..
  • Interestingly, the alignment scores between Llama2-70B and Llama3-70B are comparable to those between Llama3-70B and T5-11B (refer last rows of table 3 and table 4). There are minimal architectural differences between Llama2 and Llama3-70B (primarily in the context window size). This implies that factors such as training regimes (e.g., dataset size and composition) and context window size are as influential as model architecture in achieving strong alignment.
  • Table 4 shows the cross-architecture alignment scores for different LLMs, comparing models of similar sizes but varying architectures. Despite architectural differences, models of similar sizes exhibit moderate to high alignment scores (mostly above 0.5), suggesting a consistent semantic organization in LLMs.
Overall, we observed moderate to high alignment across LLMs, regardless of their size, architecture, or pretraining, indicating that the findings in Section 4 may generalize to other LLMs.
6 Bias Mitigation: Case Study of Cluster Modification
To investigate RQ4 and demonstrate the practical impact of our findings, we conduct a case study on modifying conceptual clusters to mitigate ethnicity bias. A key challenge here is balancing bias reduction with the preservation of the model’s overall utility and linguistic integrity. Our approach focuses on embedding engineering, targeting clusters of tokens associated with stereotype-prone identities, such as proper nouns from the Indian subcontinent (e.g., Indian and Pakistani human names). We hypothesize that by modifying these token embeddings, we can disrupt learned biases without sacrificing the tokens’ semantic roles within the language model. For example, we aim to investigate if the token “Sharma”, a member of the Indian proper nouns cluster, retains its identity as a proper noun after token engineering. To evaluate fairness score, we use Bias Benchmark for Question Answering (BBQ) dataset 37, the fairness evaluation methodology is based on the WinoBias evaluation59. To assess the linguistic preservation, we chose the part-of-speech (POS) tagging task as a proxy indicator on CoNLL-2003 dataset 52 as well as a subset of Wikimedia dataset 58 (Further evaluation on the robustness of embedding engineering on GLUE, SuperGLUE, and SQUAD 55, 54 benchmarks is provided in Appendix D to provide a more comprehensive assessment of general language model quality after the token manipulations).
Our approach begins by selecting a cluster of token embeddings from the original model. For this experiment, we used the communities associated with Indian and Pakistani human names that was identified by our concept extraction algorithm. Then, we calculated the mean and standard deviation of the joint cluster. Finally we modified these clusters to form a single cluster with Gaussian distribution of the joint cluster (i.e. samples from the $Gaussian(\mu,0.7*\sigma)$). Then, we finetuned the model for POS tagging task for CoNLL-2003 dataset fro 5 epochs (see Appendix B for more details). For the POS tagging task, we utilize the CoNLL-2003 dataset 52, a widely-used benchmark for named entity recognition and POS tagging. Each model is fine-tuned using a standard supervised fine-tuning approach.
As shown in Table 5, all token-engineered models exhibit over 90% token overlap with their base counterparts. This high overlap (like the >90% reported) suggests that even though the embeddings have been modified, they still largely represent the same underlying POS tags. Furthermore, for those POS tags where the base and modified models show disagreement, we still observe the same overall quality. This means that even where the model does change its POS tag assignment, the new assignment is just as likely to be correct as the original. At the same time, we are able to mitigate bias ranging from 44% for the Albert-xxl-large and T5-3b models to 52% for the T5-11b model. These highlight the potential of token engineering for bias mitigation while keeping the general purpose utility of the model intact.
Table 5: Fairness and POS Tagging Performance of Various Models. Fairness: higher score implies less bias. POS: higher score implies better accuracy.
Fairness POS Tagging
Model Base Ours Base Our Overlap
Score Score Acc Acc %
Albert-base 0.26 0.74 0.91 0.91 0.90
Albert-xxl 0.28 0.72 0.93 0.91 0.92
T5-3b 0.28 0.72 0.92 0.91 0.94
T5-11b 0.24 0.76 0.94 0.93 0.94
7 Conclusion
In this paper, we propose a modular concept extraction mechanism that uncovers the emergence of distinct conceptual communities within the entire input embedding space. Using our methodology, we observe that LLMs form organized conceptual structures within their input embedding spaces. We demonstrate that the input embeddings of LLMs form categorical semantic structures that align with external world representations. We quantitatively analyze several properties of these structures, with a particular focus on categorical structures related to named entities. Additionally, we observe that numerical structures within the input embedding layer align with human notion of numerical values, including a topological ordering of numbers. We also discussed that LLMs inherently exhibit a degree of alignment with one another, suggesting the potential to extend the observed human-LLM alignment to other models. This study opens new avenues for further exploration and intervention in LLMs, especially within the realm of embedding engineering in several key areas, including bias detection and mitigation.
8 Limitations and Risks
Risks: This work provides detailed information about (1) the formed clusters/concept in the input embedding layer, and (2) the separation of memory from reasoning in Albert. As the methodology can also be applied to other models, it can potentially facilitate more advanced adversarial attacks and content manipulation in LLMs.
Appendix A Semantic Representation Space
A.1 conceptual groupings, Representation Learning and Semantic Memory
A simplified model of human understanding can be described as a multi-step process in which the incoming sensory information is integrated and turned into a brain-constructed interpretation of external objects/stimuli. This results in a mental construct (or percept) 16. Then, the newly formed mental construct gets compared, integrated, and associated with the existing knowledge already retained in the semantic memory. These connections help us to conceptualize and understand the new information based on what we already know (as shown in Figure 3). While diverse reasoning mechanisms exist, a common thread among them seems to be their reliance on and interaction with the semantic memory. This interaction likely leaves enduring traces within such a memory 5. The integration of newly encountered mental constructs within an existing cognitive framework is often guided by their relations to established internal constructs 14. This process facilitates the creation, refinement, or expansion of semantically similar clusters 28, implying a degree of inherent categorization [^3] This notion implies an interrelation between the recognition and categorization..
Representation Learning and Conceptual roupings) Representation learning is a fundamental aspect of language models, where the goal is to learn distributed representations (embeddings) for words or subword units that capture their semantic and syntactic relationships 4. Early models like word2vec 33 utilized shallow neural networks to learn these embeddings from large text corpora. More recent models, such as Large Language Models (LLMs) like GPT-3 30, employ transformer architectures with self-attention mechanisms 53, enabling more accurate and dynamic representation learning. From the distributional hypothesis, models can form "concepts" by identifying patterns and relationships within data, particularly through recognizing the approximate invariance of shared features across different data instances 14. It should be noted that "conceptual groupings" is a byproduct of the learning process, not a guaranteed outcome. Furthermore, the formed "concepts" might not always align with human-defined concepts. Thus, it is crucial to investigate and evaluate the nature of conceptual groupings within models and their alignment with human understanding.
A.2 Evaluation Methods
Intrinsic evaluation methods, assess semantic representations by measuring word embedding similarity, comparing them to human judgments of relatedness, (e.g. 33, 18, 51, 35, 49). While these methods provide valuable insights, they often suffer from the limitations that probing techniques impose, and may not fully capture how well a model forms conceptual hierarchies or grasps the full spectrum of semantic relationships, as word embeddings typically represent individual words rather than higher-level categories 10. This limitation can hinder the evaluation of a model’s ability to understand broader semantic relationships and its capacity for abstract reasoning.
Another approach is to use extrinsic evaluation methods, which measure the performance of language models on downstream tasks that rely on semantic understanding, such as sentiment analysis, question answering, and machine translation 55, 54. The performance on these tasks can indirectly indicate the quality of the learned semantic representations. The main challenge with using such extrinsic evaluation metrics to assess conceptual groupings in the embedding space is that they do not directly measure the quality of the concepts themselves. Instead, they measure how well the model performs on specific tasks that rely on those concepts, which does not necessarily prove that the model has formed robust, human-like concepts. For instance, a model might accurately classify sentiment without truly grasping the nuances of emotions like sarcasm or irony 31.
In addition to evaluation metrics, researchers have also employed visualization methods to interpret learned representations. These methods often involve projecting high-dimensional embeddings into lower-dimensional spaces for visualization 46, 44. However, several challenges arise with these approaches. Firstly, they offer indirect and subjective assessments of model understanding, lacking a quantitative basis for evaluation. Secondly, the dimensionality reduction process can lead to information loss 15, 56, potentially obscuring important nuances in the representations. Finally, different visualization techniques can produce conflicting results, making it difficult to reach definitive conclusions about the model’s true comprehension. To mitigate these shortcomings, we propose a concept extraction mechanism that identifies communities in higher-dimensional space. This approach enables both quantitative evaluation and mitigation of potential information loss incurred during dimensionality reduction.
Appendix B Bias Mitigation details
We evaluated the models on subsets of the Wikimedia dataset 58, where we sample only those sentences that contain the tokens from our token-engineered cluster. For the fairness task, we adopt the Bias Benchmark for Question Answering (BBQ) dataset 37, focusing on the nationality split. We modify the data set for masked language modeling (MLM) by replacing the interrogative questions with a token ’[MASK]’. To ensure the quality of evaluation, each sample is manually checked for grammatical correctness following this transformation. For evaluating fairness, we compare each base model with its corresponding token-engineered model that utilizes Gaussian-sampled embeddings. We first filter out the evaluation samples that pass the fairness test on the base model, thus isolating only the problematic cases. Our evaluation of fairness is inspired by the evaluation metrics used for the winobias dataset 59.For the remaining biased samples, we calculate which model (base or token-engineered) is more likely to generate a biased output by examining their output logit probability.
Appendix C Lovain Algorithm
@@ 15       Reduce each community to a single node;
Evaluation Metrics. conceptual groupings involves the creation of abstract internal representations through the clustering of inputs with shared features. While numerous intrinsic clustering metrics exist to assess cluster formation quality, their application to our use case is limited by two factors:
@@
  • The formed clusters are not situated within Euclidean space 24, 9, 8, rendering geometric properties such as cluster distances inadequate indicators of concept well-formedness and distinctiveness.
@@
  • Concepts inherently possess a degree of vagueness 20, thus metrics like compactness or separation do not reliably reflect the quality of formed concepts.
Appendix D Knowledge-Reasoning Separation
Table 6: Huggingface Albert base model on GLUE, SuperGLUE, and SQUAD tasks. For the baseline, the model was finetuned without altering the embeddings. For the mid-point, the embedding layer entries are assigned the mid-point embedding of their associated community. We used 41 repo for benchmarking.
mid-point GLUE
Method/Tasks mnli mrpc qnli qqp rte sst stsb wnli
baseline 0.827 0.841 0.902 0.858 0.765 0.915 0.872 0.549
mid-point 0.849 0.865 0.910 0.875 0.783 0.922 0.890 0.563
mid-point SuperGLUE SQUAD SQUAD
Method/Tasks boolq cb copa multirc record wic wsc v1 (f1) v2 (f1)
baseline 0.622 0.512 0.59 0.350 0.586 0.595 0.528 83.72 70.9
mid-point 0.621 0.478 0.55 0.372 0.588 0.626 0.634 84.4 74.9
Now, we are interested to see the extent to which the knowledge learned during pretraining is modular and separable from the reasoning mechanisms employed by the language model in downstream tasks. For example, can the knowledge learned during the pretraining phase be selectively removed or modified without significantly impacting the model’s performance on finetuning? The modularity can also impact the effectiveness of Language Model Inversion 34, 47 techniques, which aim to extract private information such as names or other sensitive details learned during the pre-training444 For example, if knowledge is found to be highly modular, it may be possible to develop targeted interventions that obscure or remove specific sensitive information without significantly impacting the model’s overall performance on downstream tasks..
Appendix E Hierarchical Community Extraction: Methodology and Considerations
There are multiple viable strategies to extract hierarchical communities in our methodology; the first one is to use algorithms such as Louvain that inherently generate hierarchical communities 6. However, the summation of the weights as a new weight in the community aggregation phase of Louvain algorithm skews the weighted graph in favor of merging smaller communities in the next phase. This detaches the community detection from the actual values in the semantic representation space (i.e. graph weights higher in the Louvain hierarchy no longer reflect the geometrical affinity of the nodes). Thus, for our concept extraction algorithm, we use 1-2 community aggregation and rather use KNN iteratively with different granularity for extracting hierarchical concepts (more details are described in Algorithm 1).
Note that we only use well-established methods such as $K$-NN, UMAP-based weighing formula, as well as Louvain 6 and label propagation 60 community detection algorithms to capitalize on the established generalizability of these algorithms. Although, as we show in the next sections that our method produces amazingly good categories, it should be noted that we intend to focus our analysis on “if the language model forms concepts" rather than creating the most optimal concept extraction mechanism. Thus, as an extension to this work, one can focus on further optimizing our proposed method. Notably, our concept extraction is algorithm-agnostic; alternatives could be readily employed.
Appendix F Models Under Investigation
We used huggingface repository for all our models.
F.1 Albert
Albert (A Lite BERT) 23 is a transformer-based model for language representation learning, designed to be more efficient than its predecessor, BERT 11. While it shares the same basic architecture as BERT, it incorporates two main key modifications:
Note that reducing the number of parameters and sharing information across layers can force the model to learn more general representations, thus indirectly contributing to better conceptual groupings. datasets. Albert is pretrained on English text datasets, namely the English Wikipedia and BookCorpus, using self-supervised learning objectives.
Tokenization. It uses Sentencepiece tokenizer 22 on the uncased corpus with a vocabulary size limit of 30K tokens.
F.2 T5
F.3 GloVe
Tokenization. They used Stanford tokenizer  [^7] https://nlp.stanford.edu/software/tokenizer.shtml, a form of BPE tokenization scheme 45, 13 that constructs unigram counts from a corpus, and optionally thresholds the resulting vocabulary based on total vocabulary size (2.2M 8882196016 cased tokens most frequent words in the case of GloVe embeddings) or minimum frequency count 40.
F.4 Llama3
LLaMA 3 (Large Language Model Meta AI) is a decoder-only large language model 1, which Grouped Query Attention (GQA) that allows the model to effectively handle longer contexts. The model is available in various sizes, including 1B, 3, and 70B parameters. The larger models exhibit significantly improved capabilities in reasoning and complex language tasks.
Training Dataset. LLaMA 3 is pretrained on an extensive dataset, including over 15 trillion tokens sourced from diverse text corpora, such as books, articles, and websites. This large-scale training ensures comprehensive language understanding across different domains .
Appendix G Case Sensitivity Analysis